OpenStudy (anonymous):

given log6 + xlog4 = log4 +log(32+4^x) find x

OpenStudy (anonymous):

\[\log a+\log b=\log(a\times b)\]

OpenStudy (anonymous):

\[\log a-\log b=\log(\frac{a}{b})\]

OpenStudy (anonymous):

\[\large n\log x=\log x^n\]

OpenStudy (anonymous):

thanks a lot :)